# AB-UPT: Scaling Neural CFD Surrogates for High-Fidelity Automotive Aerodynamics Simulations via Anchored-Branched Universal Physics Transformers

### v3

@@ 3.3 Enforcing physics consistency
As discussed in Section 2.1, physical fields are often intrinsically governed by local differential operators, enforcing constraints that reflect fundamental conservation laws. However, imposing differential-operator constraints on neural networks requires a field-based formulation, allowing pointwise derivative computation. Conventional Transformer architectures cannot be directly viewed as continuous fields because they map input sets to output sets. By contrast, as described in Section 3.2, our anchored neural field architecture treats the network as a conditional field, using anchor tokens’ key/value vectors as the context to produce pointwise predictions at arbitrary query positions. This approach enables the application of local differential operators. To demonstrate and leverage the potential of the anchored-field formulation, we model the vorticity as an intrinsically divergence-free field, constructed from the local rotation of the velocity (like the real physical field), thus ensuring physical consistency.
@@
A well-known mathematical property of the vorticity (and more generally the curl of a field) is that its divergence vanishes, i. e., (×$\nabla\cdot(\nabla\times\bm{u})=0$. Intuitively, this means that vorticity has no monopole-like sources or sinks: vortex lines can only form closed loops or extend to infinity, never beginning or ending at a point.
@@
We found that computing the spatial derivatives in Equation 10 via PyTorch’s automatic differentiation is inefficient due to suboptimal support of vectorized cross-attention in torch.compile. We therefore instead compute the spatial derivatives with a central-difference approximation,
@@
where $\hat{u}_{i}$ is the i-th (scalar) velocity output, $\bm{e}_{j}$ is the unit vector in the $x_{j}$-direction, and $\delta$ is the step-size. Here, the dot in $\hat{\bm{u}}(\bm{x},\,\cdot)$ indicates that the prediction also depends on additional inputs such as the anchor token’s key and value vectors, which we omit for notational simplicity. To approximate the partial derivatives w. r. t. each of the three spatial coordinates, we perform six additional forward passes, each applying an offset $+\delta\bm{e}_{j}$ and $-\delta\bm{e}_{j}$ to the input position.
@@ 4 Experiments
We begin our experiments by showing that neural surrogates do not need to process the full mesh to obtain accurate models (Section 4.2), which enables efficient training on large problem-scales. Using this insight, we design Anchored-Branched Universal Physics Transformers (AB-UPT) by adapting the UPT architecture (Alkin et al., 2024a) to handle complicated physical dynamics of CFD simulations with million-scale meshes (Section 4.3). To demonstrate the effectiveness of AB-UPT, we compare it against other neural surrogate models in Section 4.4. Next, we demonstrate how AB-UPT can tackle practical use-cases such as calculating aerodynamic drag and lift coefficients (Section 4.5), simulating directly from a CAD geometry without the expensive creation of a simulation mesh (Section 4.6) as well as enforcing physical consistency via a divergence-free vorticity formulation (Section 4.7). Finally, we show scaling behavior of AB-UPT w.r.t. problem, model and data scales in Section 4.8. Additional experiments are provided in Appendix A. The experimental setup for AB-UPT is outlined in Appendix B.3
Table 2: Overview of considered datasets. Datasets vary in the number of surface/volume points (#Points) and their variables. High-quality datasets contain surface pressure $\bm{p}_{s}$ and wallshearstress $\bm{\tau}_{w}$ as well as volume pressure $\bm{p}_{v}$, velocity $\bm{u}$ and vorticity $\bm{\omega}$. As CFD simulations are expensive to compute and store, these datasets do not contain many simulations, which also need to be split into train/validation/test splits.
@@
Dataset #Points $\bm{p}_{s}$ $\bm{\tau}_{w}$Points $\bm{p}_{v}$ $\bm{u}$ $\bm{\omega}$Simulations Size Cost [^2] Cost estimated with \$3 per Amazon EC2 hpc6a.x48xlarge node-hour which was used for both AhmedML and DrivAerML. This is the publicly visible on-demand hourly rate. Actual costs may vary based instance provisioning model.
@@ 4.2 Efficient training on million-scale CFD meshes
Figure 5: Train and test losses with varying number of training inputs. Contrary to numerical solvers, neural surrogates do not require million-scale meshes to train accurate models. Training on more points beyond a dataset-specific threshold results in overfitting and requires larger datasets (either in terms of data-scale, i.e., number of simulations or problem-scale, i.e., complexity of the problem) to benefit from the increased modeling capacity when using more points, which is costly. This can be observed in the first and second row where AhmedML and DrivAerML have the same data scale (400 train simulations) but AhmedML has smaller problem scale, resulting in overfitting. Note that increasing data- or model-scale is orthogonal to our work. Although runtime increases with a larger number of inputs, the amount of GPU-hours required is still low, and scaling to even larger problem-scales could be easily implemented via sequence parallelism (Liu et al., 2023), which allows scaling quadratic self-attention to millions of tokens. Note that Transformer/AB-UPT requires less memory due to constant memory complexity of FlashAttention (Dao et al., 2022), whereas Transolver uses memory heavy torch.einsum operations to create slices. See Appendix A.7 for more details on Transolver scaling.
@@
We show in Figure 5 that it is feasible to train standard Transformer self-attention to accurately predict simulation results on million-scale meshes by training on lower resolutions, outperforming linear transformer architectures (such as Transolver). As self-attention requires quadratically more compute w. r. t. the number of inputs, computational demands increase. However, the number of GPU-hours required to train models is still low (models with 16 thousand surface/volume points train in less than 10 GPU-hours). Given even more complicated physical simulations, it would be no issue to scale our AB-UPT to more GPUs via sequence parallelism (Liu et al., 2023), which distributes tokens among multiple GPUs, to allow processing millions of tokens despite the quadratic complexity of self-attention. Due to the low amount of GPU-hours required, even for physical dynamics that arise from meshes with over 100 million cells, AB-UPT could be easily scaled to even more complicated problems. For reference, foundation models in domains such as computer vision (e. g., Oquab et al. (2023)) use 10 to 100 thousand GPU-hours, and compute requirements for language models can be millions of GPU-hours (e. g., Dubey et al. (2024)).
Note that random subsampling, as we apply it to create a low-resolution input mesh, is also a form of data augmentation where a network will never see the same constellation of inputs twice during training. As high-quality numerical CFD simulations are expensive to obtain (e. g., DrivAerML costs \$1M, see Table 2), datasets often have less than a thousand samples, where this form of regularization helps to prevent overfitting.
@@ 4.4 Benchmarking AB-UPT against other neural surrogate models
PointNet 12.01 3.05 8.51 5.43 67.29 23.26 28.12 880.0
GRAPH U-NET 10.33 2.48 6.42 4.14 53.53 16.13 17.98 330.0
GINO 13.27 2.52 7.90 6.23 71.80 13.03 40.58 98.5
LNO 9.05 2.29 12.91 7.56 72.04 20.51 23.27 173.0
@@
In this section, we compare the performance of AB-UPT against established neural-surrogate models: Table 4 presents the relative L2 error for surface pressure, volume velocity, and vorticity. We compare AB-UPT against the following models: PointNet (Qi et al., 2017a), Graph U-Net (Gao & Ji, 2019), GINO (Li et al., 2023b), LNO (Wang & Wang, 2024), UPT (Alkin et al., 2024a), OFormer (Li et al., 2023a), Transolver (Wu et al., 2024), Transformer (Vaswani et al., 2017). We do not consider Transolver++ as a baseline due to reproducibility issues, which we explain in Appendix B.4 and C.2. These comparisons are conducted across three diverse datasets: ShapeNet-Car (Umetani & Bickel, 2018), AhmedML (Ashton et al., 2024a), and DrivAerML (Ashton et al., 2024b).
@@
The 16K/16K surface/volume points for AhmedML and DrivAerML are motivated by the insights of Figure 5, showing that 16K obtains good results at relatively low compute costs. Evaluation procedure and evaluation metrics are detailed in Appendix B. Appendix B also presents the training hyperparameters and the number of model parameters for each model. Additionally, Appendix C.1 provides a statement on the fairness of our experimental setup in comparison to (Wu et al., 2024). Comprehensive results, including wall shear stress and volume pressure, are reported in Appendix A.4. Additional results for varying the number of slices in Transolver is provided in Appendix A.7 and comparison against Erwin (Zhdanov et al., 2025) is conducted in Appendix A.5.
@@
Table 4 shows that AB-UPT consistently outperforms all other neural surrogate models across every evaluated metric and dataset, often by a substantial margin. Moreover, we observe that most baselines struggle with properly modeling the complex vorticity. AB-UPT, however, maintains strong performance for both surface and volume metrics, highlighting its robustness. For most metrics, either a plain Transformer or Transolver is the second-best-performing model. However, due to the full self-attention mechanism of the Transformer, this model will not be able to scale to industry-standard simulation meshes (Transformer corresponds to the third row of Table 3, which shows roughly 100 times slower test speed than AB-UPT). Transolver, on the other hand, would require sequence parallelism over multiple GPUs to handle large simulation meshes (as has been explored in (Luo et al., 2025)). We can conclude that AB-UPT is the best performing model for neural-surrogate automotive CFD simulation, while also having favorable properties concerning problem scaling.
@@ 4.6 Training from CAD
Table 5 compares the accuracy of such approaches. Sampling anchor positions from the CFD simulation mesh implicitly assigns more computational resources to highly resolved regions, which typically contain complicated dynamics. Uniformly sampling from the CAD geometry removes this bias and degrades performance slightly, as expected. However, this performance degradation is fairly small, particularly when considering that numerical simulations would produce completely unusable results.
@@
AB-UPT 3.70 2.50 7.59 4.14 6.64 76.5 Geometry mesh (CAD) Regular grid
@@ 4.7 Divergence-free vorticity
As discussed in Sections 3.2 and 3.3, the conditional field property of AB-UPT allows for defining vorticity field predictions that are divergence-free by design. We train the model in two stages: in the pretraining phase, we learn a model that predicts all surface and volume fields except the vorticity, using the same model configuration and hyperparameters as in the previous experiments; in the finetuning phase, we learn all fields including the derived vorticity field (cf. Section 3.3). Furthermore, we use float32 precision for finetuning, and we train with a smaller initial learning rate and fewer training steps: starting from $2e^{-5}$, the learning rate is decayed to $1e^{-6}$ via the cosine annealing schedule. As discussed in Section 3.3, after correcting for the scaling of the velocity field due to normalization, the predicted vorticity is already in physics space. To bring the predictions to a similar scale as other predicted fields, we transform both the predictions and targets as follows: first divide by the field’s standard deviation; then apply a sign-preserving square-root to the vector’s magnitude. This is equivalent to scaling the individual loss terms. All other fields are simply normalized as in all other experiments, subtracting the mean and dividing by the standard deviation (c. f. Section B.4.1).
Table 6 shows that the divergence-free vorticity formulation of AB-UPT matches the performance of the direct-prediction approach. However, while direct prediction results in non-zero divergence of the predicted vorticity field, the divergence-free formulation guarantees zero divergence by construction. Note that the most competitive baselines such as Transolver and Transformer do not have a (conditional) field formulation, and, thus, the local divergence operator is not defined for these models.
Table 6: Evaluation of AB-UPT with the divergence-free vorticity formulation (c. f. Section 3.3). For each dataset, the first 3 columns per datasets show the relative L2/L1 test errors (in %) of surface pressure $\bm{p}_{s}$, volume velocity $\bm{u}$ and volume vorticity $\bm{\omega}$. The last column shows the mean absolute value of the divergence of the predicted vorticity field $\nabla\cdot\bm{\omega}$ (in $m^{-1}s^{-1}$), where the average is taken over simulations and samples of the field.
@@ 4.8 Scaling AB-UPT
Figure 7: Left and middle: Scaling the number of anchor tokens during inference only. Models are trained with varying anchor tokens (as indicated by their starting point and label). These models are then evaluated with an increasing number of anchor tokens but without any retraining (as indicated by the x-axis). Different colors correspond to different numbers of anchor tokens used during training. Increased number of anchor tokens used in training yields performance improvements; increased number of anchor tokens during inference yields similar accuracies. Right: Summary of the left and middle plot where the number of inference anchors is fixed at 262K (i.e., the best obtained loss) and the x-axis varies the number of training anchors instead. The performance saturation hints at the problem scale of the DrivAerML dataset, i.e., certain amounts of anchor tokens are enough to capture the intricacies of the dataset.
@@
We train AB-UPT models with various numbers of anchor tokens (4K to 131K) and additionally increase the number of anchor tokens further during inference only (up until 262K), without any retraining. By increasing the number of anchor tokens, we increase coverage of the 150M simulation mesh cells and observe saturating effects when increasing the number of anchor tokens beyond a certain point. Figure 7 shows results thereof. The left and middle plots show that AB-UPT benefits from an increased number of anchor tokens in inference time, but the best models are obtained by also training with more anchor tokens. As the number of anchor tokens increases, performance gains saturate (e.g., doubling training anchors from 4K to 8K decreases the test loss by 8%, doubling anchors from 32K to 64K decreases test loss by only 2%).

### v4

@@ 3.3 Enforcing physics consistency
As discussed in Section˜2.1, physical fields are often intrinsically governed by local differential operators, enforcing constraints that reflect fundamental conservation laws. However, imposing differential-operator constraints on neural networks requires a field-based formulation, allowing pointwise derivative computation. Conventional Transformer architectures cannot be directly viewed as continuous fields because they map input sets to output sets. By contrast, as described in Section˜3.2, our anchored neural field architecture treats the network as a conditional field, using anchor tokens’ key/value vectors as the context to produce pointwise predictions at arbitrary query positions. This approach enables the application of local differential operators. To demonstrate and leverage the potential of the anchored-field formulation, we model the vorticity as an intrinsically divergence-free field, constructed from the local rotation of the velocity (like the real physical field), thus ensuring physical consistency.
@@
A well-known mathematical property of the vorticity (and more generally the curl of a field) is that its divergence vanishes, i. e., $\nabla\cdot(\nabla\times\bm{u})=0$. Intuitively, this means that vorticity has no monopole-like sources or sinks: vortex lines can only form closed loops or extend to infinity, never beginning or ending at a point.
@@
We found that computing the spatial derivatives in Equation˜10 via PyTorch’s automatic differentiation is inefficient due to suboptimal support of vectorized cross-attention in torch.compile. We therefore instead compute the spatial derivatives with a central-difference approximation,
@@
where $\hat{u}_{i}$ is the i-th (scalar) velocity output, $\bm{e}_{j}$ is the unit vector in the $x_{j}$-direction, and $\delta$ is the step-size. Here, the dot in $\hat{\bm{u}}(\bm{x},\,\cdot)$ indicates that the prediction also depends on additional inputs such as the anchor token’s key and value vectors, which we omit for notational simplicity. To approximate the partial derivatives w. r. t. each of the three spatial coordinates, we perform six additional forward passes, each applying an offset $\delta\bm{e}_{j}$ and $\delta\bm{e}_{j}$ to the input position.
@@ 4 Experiments
We begin our experiments by showing that neural surrogates do not need to process the full mesh to obtain accurate models (Section 4.2), which enables efficient training on large problem-scales. Using this insight, we design Anchored-Branched Universal Physics Transformers (AB-UPT) by adapting the UPT architecture (Alkin et al., 2024a) to handle complicated physical dynamics of CFD simulations with million-scale meshes (Section 4.3). To demonstrate the effectiveness of AB-UPT, we compare it against other neural surrogate models in Section 4.4. Next, we demonstrate how AB-UPT can tackle practical use-cases such as calculating aerodynamic drag and lift coefficients (Section 4.5), simulating directly from a CAD geometry at inference time without the expensive creation of a simulation mesh (Section 4.6) as well as enforcing physical consistency via a divergence-free vorticity formulation (Section 4.7). Finally, we show scaling behavior of AB-UPT w.r.t. problem, model, and data scales in Section 4.8. Additional model design ablations and benchmarks are provided in Appendix A and B. The experimental setup for AB-UPT is outlined in Appendix C.3. The inference code for AB-UPT is available at: https://github.com/Emmi-AI/AB-UPT.
Unless stated otherwise, we train AB-UPT in fp16 mixed precision, circumventing previously reported instabilities (Alkin et al., 2024a) via fp32 positional embeddings and RoPE (Su et al., 2024). See Appendix A.6 for more details.
Table 2: Overview of considered datasets. Datasets vary in the number of surface/volume points (#Points) and their variables. High-quality datasets contain surface pressure $\bm{p}_{s}$ and wallshearstress $\bm{\tau}$ as well as volume pressure $\bm{p}_{v}$, velocity $\bm{u}$ and vorticity $\bm{\omega}$. As CFD simulations are expensive to compute and store, these datasets do not contain many simulations, which also need to be split into train/validation/test splits.
@@
Dataset #Points $\bm{p}_{s}$ $\bm{\tau}$ #Points $\bm{p}_{v}$ $\bm{u}$ $\bm{\omega}$ #Simulations Size [^2] Size denotes the total size of the data necessary for training AB-UPT, which only includes positions and the predicted surface/volume variables. The full raw dataset would be even larger (e.g., $\sim$ 30TB for DrivAerML). Cost [^3] Cost estimated with \$3 per Amazon EC2 hpc6a.x48xlarge node-hour which was used for both AhmedML and DrivAerML. This is the publicly visible on-demand hourly rate. Actual costs may vary based instance provisioning model.
@@ 4.2 Efficient training on million-scale CFD meshes
Figure 5: Train and test losses with varying number of training inputs. Contrary to numerical solvers, neural surrogates do not require million-scale meshes to train accurate models. Training on more points beyond a dataset-specific threshold results in overfitting and requires larger datasets (either in terms of data-scale, i.e., number of simulations, or problem-scale, i.e., complexity of the problem) to benefit from the increased modeling capacity when using more points, which is costly. This can be observed in the first and second row, where AhmedML and DrivAerML have the same data scale (400 train simulations), but AhmedML has a smaller problem scale, resulting in overfitting. Note that increasing data- or model-scale is orthogonal to our work. Although runtime increases with a larger number of inputs, the amount of GPU-hours required is still low, and scaling to even larger problem-scales could be easily implemented via sequence parallelism (Liu et al., 2023), which allows scaling quadratic self-attention to millions of tokens. Note that Transformer/AB-UPT requires less memory due to constant memory complexity of FlashAttention (Dao et al., 2022), whereas Transolver uses memory-heavy torch.einsum operations to create slices. See Appendix B.5 for more details on Transolver scaling.
@@
We show in Figure˜5 that it is feasible to train standard Transformer self-attention to accurately predict simulation results on million-scale meshes by training on lower resolutions, outperforming linear Transformer architectures (such as Transolver). As self-attention requires quadratically more compute w. r. t. the number of inputs, computational demands increase. However, the number of GPU-hours required to train models is still low (models with 16 thousand surface/volume points train in less than 10 GPU-hours). Given even more complicated physical simulations, it would be no issue to scale our AB-UPT to more GPUs via sequence parallelism (Liu et al., 2023), which distributes tokens among multiple GPUs, to allow processing millions of tokens despite the quadratic complexity of self-attention. Due to the low amount of GPU-hours required, even for physical dynamics that arise from meshes with over 100 million cells, AB-UPT could be easily scaled to even more complicated problems. For reference, foundation models in domains such as computer vision (e. g., Oquab et al. (2023)) use 10 to 100 thousand GPU-hours, and compute requirements for language models can be millions of GPU-hours (e. g., Dubey et al. (2024)).
Note that random subsampling, as we apply it to create a low-resolution input mesh, is also a form of data augmentation where a network will never see the same constellation of inputs twice during training. As high-quality numerical CFD simulations are expensive to obtain (e. g., DrivAerML costs \$1M, see Table˜2), datasets often have less than a thousand samples, where this form of regularization helps to prevent overfitting.
@@ 4.4 Benchmarking AB-UPT against other neural surrogate models
PointNet 12.09 3.05 8.02 5.44 66.04 23.63 28.13 1747.7
GRAPH U-NET 10.33 2.49 6.46 4.15 53.66 16.13 17.98 540.6
GINO 13.28 2.53 7.90 6.23 71.81 13.03 40.58 131.7
LNO 9.05 2.29 12.95 7.59 72.49 20.51 23.27 493.8
@@
In this section, we compare the performance of AB-UPT against established neural-surrogate models: Table 4 presents the relative L2 error for surface pressure, volume velocity, and vorticity. We compare AB-UPT against the following models: PointNet (Qi et al., 2017a), Graph U-Net (Gao & Ji, 2019), GINO (Li et al., 2023b), LNO (Wang & Wang, 2024), UPT (Alkin et al., 2024a), OFormer (Li et al., 2023a), Transolver (Wu et al., 2024), Transformer (Vaswani et al., 2017). We do not consider Transolver++ as a baseline due to reproducibility issues, which we explain in Appendix C.4 and B.7. These comparisons are conducted across three diverse datasets: ShapeNet-Car (Umetani & Bickel, 2018), AhmedML (Ashton et al., 2024a), and DrivAerML (Ashton et al., 2024b). A comparison on DrivAerNet++ (Elrefaie et al., 2024b) is conducted in Appendix B.1.
@@
The 16K/16K surface/volume points for AhmedML and DrivAerML are motivated by the insights of Figure 5, showing that 16K obtains good results at relatively low compute costs. Evaluation procedure and evaluation metrics are detailed in Appendix C. Appendix C also presents the training hyperparameters and the number of model parameters for each model. Additionally, Appendix B.6 provides a statement on the fairness of our experimental setup in comparison to (Wu et al., 2024). Comprehensive results, including wall shear stress and volume pressure, are reported in Appendix B.8. Additional results for varying the number of slices in Transolver is provided in Appendix B.5 and comparison against Erwin (Zhdanov et al., 2025) is conducted in Appendix B.4.
@@
Table˜4 shows that AB-UPT consistently outperforms all other neural surrogate models across every evaluated metric and dataset, often by a substantial margin. Moreover, we observe that most baselines struggle with properly modeling the complex vorticity. AB-UPT, however, maintains strong performance for both surface and volume metrics, highlighting its robustness. For most metrics, either a plain Transformer or Transolver is the second-best-performing model. However, due to the full self-attention mechanism of the Transformer, this model will not be able to scale to industry-standard simulation meshes (Transformer corresponds to the third row of Table 3, which shows roughly 100 times slower test speed than AB-UPT). Transolver, on the other hand, would require sequence parallelism over multiple GPUs to handle large simulation meshes (as has been explored in (Luo et al., 2025)). We can conclude that AB-UPT is the best performing model for neural-surrogate automotive CFD simulation, while also having favorable properties concerning problem scaling.
@@ 4.6 Training from CAD
Table˜5 compares the accuracy of such approaches. Sampling anchor positions from the CFD simulation mesh implicitly assigns more computational resources to highly resolved regions, which typically contain complicated dynamics. Uniformly sampling from the CAD geometry (surface) and random positions in the domain (volume) removes this bias and degrades performance slightly, as expected. However, this performance degradation is fairly small, particularly when considering that numerical simulations would produce completely unusable results. To further decrease this performance gap, anchor tokens could be sampled from a low-fidelity CFD mesh to approximate the anchor position distribution of the high-fidelity mesh, without requiring the expensive high-fidelity mesh generation process. We leave exploration of such approaches to future work.
@@
AB-UPT 3.70 2.50 7.59 4.14 6.64 51.2 Geometry mesh (CAD) Regular grid
@@ 4.7 Divergence-free vorticity
As discussed in Sections˜3.2 and 3.3, the conditional field property of AB-UPT allows for defining vorticity field predictions that are divergence-free by design. We train the model in two stages: in the pretraining phase, we learn a model that predicts all surface and volume fields except the vorticity, using the same model configuration and hyperparameters as in the previous experiments; in the finetuning phase, we learn all fields including the derived vorticity field (cf. Section˜3.3). Furthermore, we use float32 precision for finetuning, and we train with a smaller initial learning rate and fewer training steps: starting from $2e^{-5}$, the learning rate is decayed to $1e^{-6}$ via the cosine annealing schedule. As discussed in Section˜3.3, after correcting for the scaling of the velocity field due to normalization, the predicted vorticity is already in physics space. To bring the predictions to a similar scale as other predicted fields, we transform both the predictions and targets as follows: first, divide by the field’s standard deviation; then apply a sign-preserving square root to the vector’s magnitude. This is equivalent to scaling the individual loss terms. All other fields are simply normalized as in all other experiments, subtracting the mean and dividing by the standard deviation (c. f. Section˜C.4.1).
Table˜6 shows that the divergence-free vorticity formulation of AB-UPT matches the performance of the direct-prediction approach. However, while direct prediction results in non-zero divergence of the predicted vorticity field, the divergence-free formulation guarantees zero divergence by construction. Note that the most competitive baselines, such as Transolver and Transformer, do not have a (conditional) field formulation, and, thus, the local divergence operator is not defined for these models.
Table 6: Evaluation of AB-UPT with the divergence-free vorticity formulation (c. f. Section˜3.3). For each dataset, the first 3 columns show the relative L2/L1 test errors (in %) of surface pressure $\bm{p}_{s}$, volume velocity $\bm{u}$, and volume vorticity $\bm{\omega}$. The last column shows the mean absolute value of the divergence of the predicted vorticity field $\nabla\cdot\bm{\omega}$ (in $m^{-1}s^{-1}$), where the average is taken over simulations and samples of the field.
@@ 4.8 Scaling AB-UPT
Figure 7: Left and middle: Scaling the number of anchor tokens during inference only. Models are trained with varying anchor tokens (as indicated by their starting point and label). These models are then evaluated with an increasing number of anchor tokens but without any retraining (as indicated by the x-axis). Different colors correspond to different numbers of anchor tokens used during training. An increased number of anchor tokens used in training yields performance improvements; an increased number of anchor tokens during inference yields similar accuracies. Right: Summary of the left and middle plot where the number of inference anchors is fixed at 262K (i.e., the best obtained loss) and the x-axis varies the number of training anchors instead. The performance saturation hints at the problem scale of the DrivAerML dataset, i.e., certain amounts of anchor tokens are enough to capture the intricacies of the dataset.
@@
We train AB-UPT models with various numbers of anchor tokens (4K to 131K) and additionally increase the number of anchor tokens further during inference only (up to 262K), without any retraining. By increasing the number of anchor tokens, we increase coverage of the 150M simulation mesh cells and observe saturating effects when increasing the number of anchor tokens beyond a certain point. Figure 7 shows results thereof. The left and middle plots show that AB-UPT benefits from an increased number of anchor tokens in inference time, but the best models are obtained by also training with more anchor tokens. As the number of anchor tokens increases, performance gains saturate (e.g., doubling training anchors from 4K to 8K decreases the test loss by 8%, doubling anchors from 32K to 64K decreases test loss by only 2%).
@@
Compound-scaling
We compare individually scaling axis against jointly scaling multiple axis at once. To this end, we train an AB-UPT model with 32K surface/volume anchors, model dim 384, and model depth 24.
Table 7: Relative L2 errors (in %) of surface pressure $\bm{p}_{s}$, volume velocity $\bm{u}$, volume vorticity $\bm{\omega}$ on DrivAerML for different scaling variants and a compound scaled model. Further gains could most likely be obtained by increasing #Tokens and training duration for the compound scaled model (last row), which we leave for future exploration due to computational costs.
Setting DrivAerML
#Tokens Model dim Model depth Test loss $\bm{p}_{s}$ $\bm{u}$ $\bm{\omega}$
AB-UPT 16K 192 12 0.1314 3.82 5.93 35.1
AB-UPT 128K 192 12 0.1300 3.61 5.69 32.7
AB-UPT 16K 384 12 0.1214 3.55 5.20 25.4
AB-UPT 16K 192 60 0.1212 3.47 5.20 21.0
AB-UPT 32K 384 24 0.1169 3.38 4.87 18.8